# ChronoMagic-Bench: A Benchmark for Metamorphic Evaluation of Text-to-Time-lapse Video Generation

### v1

@@ 5.1 Evaluation Models
We select ten open-source T2V models for evaluation, including both relatively advanced U-Net based models and emerging DiT-based models. All inference parameters follow the official implementation. More details about the experiment can be found in Appendix D.
@@ 5.2 Evaluation Setups
Evaluation Criteria. We assess video quality primarily from the following four aspects: (a) Visual Quality, measures the clarity, color saturation, contrast, and overall aesthetic effect, using UMT-FVD [43], an enhanced version of FVD [64]. (b) Text Relevance measures the correlation between the prompt and the video using UMTScore [43], an enhanced version of CLIPScore [23]. (c) Metamorphic Amplitude measures the diversity and dynamic changes in the video content, using the proposed Metamorphic Score. (d) Temporal Coherence measures the smoothness and logical sequence of the video content over time, using the proposed Coherence Score. Additionally, we use human evaluation to cross-verify the reliability of the four metrics.
@@ 5.3 Comprehensive Analysis
Quantitative Evaluation. We first present and analyze the results from a qualitative perspective. All input texts are from our ChronoMagic-Bench. Unlike existing benchmarks that only assess general videos, our evaluation task focuses on generating metamorphic videos, such as the construction of houses in Minecraft, the blooming of flowers, the baking of bread rolls, and the melting of ice cubes. As shown in Figure 5, almost all U-Net-based and DiT-based models are limited to generating general videos and fail to follow prompts to produce videos with significant motion and temporal spans, except for MagicTime [79] (training data contains time-lapse videos), which underscores the importance of ChronoMagic-Pro dataset. Since T2V-Zero [28] is a zero-shot video generation model, its coherence is significantly lacking, although its visual quality is acceptable. Additionally, videos generated by AnimateDiff [22] have the best visual quality and text relevance, with high clarity and accurate adherence to the prompt’s instructions. Among the emerging DiT-based video models, OpenSora 1.2 [86] stands out as a representative model that matches the performance of U-Net based methods. It is followed by OpenSoraPlan v1.1 [40], OpenSora 1.1 [86], and Latte [45]. However, due to the inherent limitations of OpenSora 1.2’s Video VAE and 2+1D DiT architecture, the generated videos are prone to flickering, particularly during significant changes, resulting in the lowest temporal coherence (CHScore). All videos generated on ChronoMagic-Bench are available on our homepage.
@@
Qualitative Evaluation. Next, we present and analyze the results of different T2V models from a qualitative perspective. Due to the lack of metrics, we propose the first MTScore and CHScore to evaluate video motion extent and coherence, with results shown in Table 3. Consistent with Figure 5, MagicTime [79], as the only model capable of generating metamorphic videos, has the highest MTScore and GPT4o-MTScore among all models. The other models, trained only on general videos, produce videos with limited motion range due to the minimal physical knowledge encoded in the models. It can also be seen that the results of the MTScore based on feature space with lower overhead and the GPT4o-MTScore based on question answering with higher overhead are roughly similar, proving the effectiveness of the proposed indicators. Additionally, ZeroScope [62] has limited metamorphic amplitude but the best coherence, while the zero-shot algorithm T2V-Zero [28] has the lowest CHScore. U-Net based and DiT-based algorithms have similar CHScore, but the former shows superior average metamorphic amplitude. For visual quality and text relevance, the emerging OpenSoraPlan v1.1 [40] and OpenSora 1.1 & 1.2 [86] have visual quality comparable to U-Net based methods, but slightly inferior text relevance and temporal coherence. Additionally, OpenSora 1.2 [86] has the lowest UMT-FVD [43] with higher color saturation, but Quantitative Evaluation shows that LaVie [69] has the highest clarity and most accurate color reproduction. Only MagicTime [79] follows the prompt to generate a time-lapse video, but the UMTScore [43] is the lowest. We infer that the UMT-FVD [43] and UMTScore [43] are inconsistent with human perception.
@@
Human Preference. Finally, we cross-validate the effectiveness of the different metrics through Human Study. We randomly select the generated videos corresponding to 16 prompts and invited 171 participants to vote, obtaining manual evaluation results. To enhance user satisfaction, we select only five representative baseline results from which users can choose. Table 6 shows the correlation between automatic metrics and human perception. It is evident that the proposed three metrics, MTScore, CHScore, and GPT4o-MTScore, are consistent with human perception and can accurately reflect the metamorphic amplitude and temporal coherence of T2V models. Notably, although most models exhibit good coherence, they have low metamorphic amplitude. In other words, they cannot generate videos with significant physical priors, such as seed germination, egg hatching, or sunrise. This is a challenge that T2V models need to overcome in the future. Additionally, as mentioned earlier, UMTScore [43] cannot accurately measure text relevance, especially in evaluating time-lapse videos, where its Kendall and Spearman coefficients are the lowest. We infer that its feature space is not suitable for time-lapse video. Appendix D.6 provides more details about human evaluation.
5.4 Extended Analysis of Closed-Source Models
In this section, we explore the performance and limitations of closed-source models, specifically U-Net based: Gen-2 [56], Pika-2.0 [33], DiT-based: Dream Machine [44], and KeLing [32]. Given the impracticality of manually testing all 1,649 prompts in ChronoMagic-Bench, we selected two hard prompts from each of the 75 categories, resulting in ChronoMagic-Bench-150.
First, we present and analyze the results from a qualitative perspective, as shown in Figure 7. For metamorphic amplitude, most methods can only generate simple time-lapse videos, such as traffic flow; only Dream Machine [44] can generate a moderately challenging full process of night-to-day transformation; no method can generate complex changes like plant growth or building construction. In terms of temporal coherence, the performance of various closed-source models is comparable, with minor visible differences. Regarding visual quality, the DiT-based methods Dream Machine [44] and KeLing [32] outperform those based on U-Net, producing more realistic plants, more accurately saturated sky colors, and clearer traffic flow. In terms of text relevance, all methods adhere to the prompt’s instructions to generate content relevant to the theme, except for Pika-2.0 [33], which mistakenly interprets day-to-night as night-to-day.
@@ 5.4 Extended Analysis of Closed-Source Models
Next, we analyze the results from a quantitative perspective, as shown in Table 4. The results are consistent with Figure 7, with Dream Machine [44] performing better in metamorphic amplitude (MTScore, GPT4o-MTScore) and Pika-2.0 [33] showing the worst text relevance (UMTScore). DiT-based methods outperform U-Net based ones in visual quality. To facilitate comparison under a unified standard, we also test open-source models on ChronoMagic-Bench-150. It is evident that for most models, the MTScore and GPT4o-MTScore are low, and they are unable to generate videos involving complex state changes. Additionally, due to the inherent limitations of UMT-FVD [43] and UMTScore [43], they fail to accurately reflect the differences between open-source and closed-source models. However, the qualitative analysis across all models demonstrates that closed-source models consistently surpass open-source models in visual quality and textual relevance. Furthermore, it is worth noting that the results within the same domain (open/closed) align with human evaluations.
| Method | Venue | Backbone | Status | UMT-FVD | UMTScore | MTScore | GPT4o-MTScore |
| Gen-2 [56] | Runway | U-Net | Close-Source | 218.99 | 2.400 | 0.373 | 2.62 |
| Pika-2.0 [33] | PikaLab | U-Net | Close-Source | 223.05 | 2.317 | 0.347 | 2.48 |
| Dream Machine [44] | LUMA | DiT | Close-Source | 214.91 | 2.387 | 0.474 | 3.11 |
| KeLing [32] | Kwai | DiT | Close-Source | 202.32 | 2.517 | 0.369 | 2.74 |
| ModelScopeT2V [66] | Arxiv’23 | U-Net | Open-Source | 230.74 | 2.783 | 0.409 | 3.01 |
| ZeroScope [62] | CVPR’23 | U-Net | Open-Source | 260.61 | 2.232 | 0.403 | 2.29 |
| T2V-zero [28] | ICCV’23 | U-Net | Open-Source | 250.22 | 2.559 | 0.399 | 2.62 |
| LaVie [69] | Arxiv’23 | U-Net | Open-Source | 210.39 | 2.714 | 0.350 | 2.50 |
| AnimateDiff [22] | ICLR’24 | U-Net | Open-Source | 239.31 | 2.837 | 0.470 | 2.62 |
| VideoCrafter2 [10] | Arxiv’24 | U-Net | Open-Source | 214.06 | 2.763 | 0.437 | 2.87 |
| MCM [80] | Arxiv’24 | U-Net | Open-Source | 244.49 | 2.282 | 0.422 | 3.06 |
| MagicTime [79] | Arxiv’24 | U-Net | Open-Source | 294.72 | 1.763 | 0.479 | 3.05 |
| Latte [45] | Arxiv’24 | DiT | Open-Source | 232.29 | 2.122 | 0.366 | 2.42 |
| OpenSora 1.1 [86] | Github’24 | DiT | Open-Source | 241.09 | 2.676 | 0.448 | 2.57 |
| OpenSora 1.2 [86] | Github’24 | DiT | Open-Source | 210.93 | 2.681 | 0.383 | 2.50 |
| OpenSoraPlan v1.1 [40] | Github’24 | DiT | Open-Source | 228.70 | 2.459 | 0.331 | 2.21 |
5.5 Guideline for T2V Models Selection
@@
With the increasing number of T2V models, the community faces challenges in selecting the most appropriate model due to the tendency of each model to showcase its best results. To address this issue, we provide a guideline according to the Table 4 for model selection based on ChronoMagic-Bench: (1) Except for MagicTime [79] and Dream Machine [44], most T2V models exhibit minimal metamorphic amplitude and cannot generate complete processes rich in physical changes, such as seed germination, sunrise, or building construction; (2) The visual quality of a single frame may be high, but when viewed in sequence, flickering often occurs, indicating poor temporal coherence. This issue is particularly evident in T2V-zero [28] and OpenSora 1.2 [86], whereas closed-source models do not exhibit this problem; (3) The emergence of Sora [8] has promoted the rapid development of DiT-based methods. Closed-source models based on DiT have comprehensively surpassed those based on U-Net. However, open-source models’ visual quality, text-following capability, and metamorphic amplitude still lag behind U-Net-based methods. We speculate that DiT-based models are more scalable and require more data, giving closed-source models a significant advantage over open-source models; (4) It is expensive to access massive data and computing resources. First, they can build datasets by crawling videos without copyright disputes. Second, adopting the U-DiT architecture may balance performance and cost to a certain extent; (5) Ordinary users who want to try T2V models can prioritize Dream Machine [44] and KeLing [32]. Researchers who wish to conduct in-depth research on T2V can prioritize the study of metamorphic video generation with OpenSoraPlan [40] and OpenSora [86], as neither open-source nor closed-source models can achieve this function.
@@ 6 Conclusion
In this paper, we present ChronoMagic-Bench, the first benchmark specifically designed to assess the generation of time-lapse videos. It addresses the shortcomings of current benchmarks, which primarily focus on standard videos and overlook critical elements such as metamorphic amplitude and coherence. Additionally, we introduce two new automated metrics, MTScore and CHScore, which align with human perception. Based on ChronoMagic-Bench, we conduct a comprehensive evaluation of almost all open-source leading text-to-video (T2V) models and provide crucial insights into the strengths and weaknesses of various models. Moreover, we propose ChronoMagic-Pro, the first large-scale time-lapse text-to-video dataset, to facilitate further research by the community.
Limitations and Future Work. While ChronoMagic-Bench offers a robust evaluation framework, there are two limitations to this work. First, although MTScore and CHScore are introduced to assess metamorphic attributes and temporal coherence, these metrics only relatively reflect the quality of different T2V models. And they do not ascertain whether the videos adhere to physical laws. For example, the MTScore of MCM [80] is relatively high, but the video changes strangely. Second, we use the best existing UMT-FVD and UMTScore to evaluate visual quality and text relevance, but due to the inapplicability of feature space, models in different domains (open/closed source) are not comparable. We will address this issue in the future.
Appendix
[chapters]1
Appendix A More Related Work: Text-to-Video Generation Models
@@ B.1 Construction of Retrieval Sentences for Metamorphic Score
B.2 More Description of Temporal Coherence Score
@@
We present a concise global description of the algorithm for computing the Temporal Coherence Score, as shown in Algorithm 1. First, we process the input video using a pre-trained model with grid size and threshold to obtain . Next, we count the number of missing tracking points in each frame and the change in missed points between consecutive frames . We then calculate the average proportion of missed points per frame , indicating the overall visibility issue across the video. Following this, we compute the variation in the number of missed points between consecutive frames , measuring frame-to-frame coherence. We also determine the ratio of frames that need to be cut , reflecting the extent of video editing required, and count the number of consecutive changes in missed points exceeding the threshold , indicating frequent large-scale instability in point tracking. Additionally, we measure the maximum continuous change in missed points , highlighting the most severe continuity breaks in the video. Finally, we integrate these metrics to calculate the Coherence Score (CHScore).
B.3 Visualization of the Different Scores of MTScore and CHScore

### v2

@@ 5.1 Evaluation Models
We select fourteen open-source T2V models for evaluation, including both relatively advanced U-Net based models (e.g., ModelScopeT2V [68], ZeroScope [64], T2V-zero [28], LaVie [71], AnimateDiff [22], MagicTime [83], VideoCrafter2 [11] and MCM [84].) and emerging DiT-based models (e.g., Latte [47], OpenSoraPlan v1.1 [41], OpenSora 1.1 [90], OpenSora 1.2 [90], EasyAnimate [77], CogVideoX [81].). We also selected four closed-source models for evaluation, specifically U-Net based: Gen-2 [58], Pika-2.0 [34], DiT-based: Dream Machine [46], and KeLing [33]. All inference settings follow the official implementation. For more details, please refer to the Appendix D.
@@ 5.2 Evaluation Setups
Evaluation Criteria. We assess video quality primarily from the following four aspects: (a) Visual Quality measures the clarity, color saturation, contrast, and overall aesthetic effect, using UMT-FVD [45], an enhanced version of FVD [66]. (b) Text Relevance measures the correlation between the prompt and the video using UMTScore [45], an enhanced version of CLIPScore [23]. (c) Metamorphic Amplitude measures the diversity and dynamic changes in the video content, using the proposed Metamorphic Score. (d) Temporal Coherence measures the smoothness and logical sequence of the video content over time, using the proposed Coherence Score. Additionally, we use human evaluation to cross-verify the reliability of the four metrics.
@@ 5.3 Comprehensive Analysis
Quantitative Evaluation. We first present and analyze the results from a qualitative perspective. All input texts are from our ChronoMagic-Bench. Unlike existing benchmarks that only assess general videos, our evaluation task focuses on generating metamorphic videos, such as the construction of houses in Minecraft, the blooming of flowers, the baking of bread rolls, and the melting of ice cubes. As shown in Figure 5, almost all U-Net-based and DiT-based models are limited to generating general videos and fail to follow prompts to produce videos with significant motion and temporal spans, except for MagicTime [83] and CogVideoX [81] (training data contains time-lapse videos), which underscores the importance of ChronoMagic-Pro dataset. Since T2V-Zero [28] is a zero-shot video generation model, its coherence is significantly lacking, although its visual quality is acceptable. Among the emerging DiT-based video models, CogVideoX [81] and OpenSora v1.2 [90] stands out as a representative that matches the performance of U-Net based methods, followed by EasyAnimate V3 [22], OpenSoraPlan v1.1 [41], while Latte [47] shows poor text-following capability.
@@
Qualitative Evaluation. Next, we present and analyze the results of different T2V models from a qualitative perspective as shown in Table 3. Consistent with Figure 5, MagicTime [83] and CogVideoX [81], as the only model capable of generating metamorphic videos, has the highest MTScore and GPT4o-MTScore among all models. The other models, trained only on general videos, produce videos with limited motion range due to the minimal physical knowledge encoded in the models. It can also be seen that the results of the MTScore based on feature space with lower overhead and the GPT4o-MTScore based on question answering with higher overhead are roughly similar, proving the effectiveness of the proposed indicators. Additionally, ZeroScope [64] has limited metamorphic amplitude but the best coherence, while the zero-shot algorithm T2V-Zero [28] has the lowest CHScore. U-Net based and DiT-based models have similar CHScore, but the former shows superior average metamorphic amplitude. For visual quality and text relevance, the emerging CogVideoX [81] and EasyAnimate [77] performed best. The OpenSoraPlan v1.1 [41] and OpenSora 1.1&1.2 [90] have visual quality comparable to U-Net based methods, but slightly inferior text relevance. Only MagicTime [83] and CogVideoX [81] follows the prompt to generate a time-lapse video, but the UMTScore [45] is the lowest. We infer that the UMT-FVD [45] and UMTScore [45] are inconsistent with human perception.
@@
Human Preference. Finally, we cross-validate the effectiveness of the different metrics through Human Study. We randomly select the generated videos corresponding to 16 prompts and invited 212 participants to vote, obtaining manual evaluation results. To enhance user satisfaction, we select only six representative baseline results and reference videos from which users can choose. Figure 6 shows the correlation between automatic metrics and human perception. It is evident that the proposed three metrics, MTScore, CHScore, and GPT4o-MTScore, are consistent with human perception and can accurately reflect the metamorphic amplitude and temporal coherence of T2V models. Additionally, as mentioned earlier, UMTScore [45] cannot accurately measure text relevance, especially in the evaluation of time-lapse videos, where its Kendall and Spearman coefficients are the lowest. We infer that its feature space is not suitable for time-lapse video. For more details please refer to Appendix D.
| Method | Venue | Backbone | Status | UMT-FVD | UMTScore | MTScore | CHScore | GPT4o-MTScore |
| Gen-2 [58] | Runway | U-Net | Close-Source | 218.99 | 2.400 | 0.373 | 125.25 | 2.62 |
| Pika-2.0 [34] | PikaLab | U-Net | Close-Source | 223.05 | 2.317 | 0.347 | 75.98 | 2.48 |
| Dream Machine [46] | LUMA | DiT | Close-Source | 214.91 | 2.387 | 0.474 | 95.97 | 3.11 |
| KeLing [33] | Kwai | DiT | Close-Source | 202.32 | 2.517 | 0.369 | 74.20 | 2.74 |
| ModelScopeT2V [68] | Arxiv’23 | U-Net | Open-Source | 230.74 | 2.783 | 0.409 | 61.01 | 3.01 |
| ZeroScope [64] | CVPR’23 | U-Net | Open-Source | 260.61 | 2.232 | 0.403 | 94.67 | 2.29 |
| T2V-zero [28] | ICCV’23 | U-Net | Open-Source | 250.22 | 2.559 | 0.399 | 18.54 | 2.62 |
| LaVie [71] | Arxiv’23 | U-Net | Open-Source | 210.39 | 2.714 | 0.350 | 81.32 | 2.50 |
| AnimateDiff V3 [22] | ICLR’24 | U-Net | Open-Source | 239.31 | 2.837 | 0.470 | 70.36 | 2.62 |
| VideoCrafter2 [11] | CVPR’23 | U-Net | Open-Source | 214.06 | 2.763 | 0.437 | 75.90 | 2.87 |
| MCM-MSLION [84] | Arxiv’24 | U-Net | Open-Source | 244.49 | 2.282 | 0.422 | 58.08 | 3.06 |
| MagicTime [83] | Arxiv’24 | U-Net | Open-Source | 294.72 | 1.763 | 0.479 | 77.98 | 3.05 |
| Latte [47] | Arxiv’24 | DiT | Open-Source | 232.29 | 2.122 | 0.366 | 72.57 | 2.42 |
| OpenSora 1.1 [90] | Github’24 | DiT | Open-Source | 241.09 | 2.676 | 0.448 | 75.94 | 2.57 |
| OpenSora 1.2 [90] | Github’24 | DiT | Open-Source | 210.93 | 2.681 | 0.383 | 51.87 | 2.50 |
| OpenSoraPlan v1.1 [41] | Github’24 | DiT | Open-Source | 228.70 | 2.459 | 0.331 | 61.50 | 2.21 |
| EasyAnimate V3 [77] | Arxiv’24 | DiT | Open-Source | 202.03 | 2.733 | 0.352 | 88.48 | 2.33 |
| CogVideoX-2B [81] | Arxiv’24 | DiT | Open-Source | 195.52 | 3.240 | 0.472 | 38.64 | 3.09 |
@@
Extended Analysis of Closed-Source Models In this section, we explore the performance and limitations of closed-source models. Given the impracticality of manually testing all 1,649 prompts in ChronoMagic-Bench, we selected two hard prompts from each of the 75 categories, resulting in ChronoMagic-Bench-150. We first analyze the results from a quantitative perspective. As shown in Table 4, with Dream Machine [46] performing better in metamorphic amplitude (MTScore, GPT4o-MTScore) and Pika-2.0 [34] showing the worst text relevance (UMTScore). DiT-based methods outperform U-Net based ones in visual quality. To facilitate comparison under a unified standard, we also test open-source models on ChronoMagic-Bench-150. It is evident that for most models, the MTScore and GPT4o-MTScore are low, and they are unable to generate videos involving complex state changes. Additionally, due to the inherent limitations of UMT-FVD [45] and UMTScore [45], they fail to accurately reflect the differences between open-source and closed-source models. However, the qualitative analysis across all models demonstrates that closed-source models consistently surpass open-source models in visual quality and textual relevance. Furthermore, it is worth noting that the results within the same domain (open/closed) align with human evaluations. We also conduct a detailed qualitative analysis, please refer to Appendix D.5 for more details.
Exploratory Experiment on ChronoMagic-Pro. To verify the validity and robustness of the ChronoMagic-Pro, we conducted quantitative validation based on OpenSoraPlan v1.1 [41]. Specifically, we fine-tuned the temporal module of the OpenSoraPlan v1.1 model using the Magic Training Strategy [83], based on the weights of OpenSoraPlan v1.1 [41]. Due to limited computational resources, we randomly selected only 10,000 video-text pairs from ChronoMagic-Pro for training. The results are shown in Table 3. After fine-tuning with ChronoMagic-Pro, the visual quality (i.e., UMT-FVD), text relevance (i.e., UMTScore), and metamorphic amplitude (i.e., MTScore and GPT4o-MTScore) were all effectively improved. Moreover, we utilized a straightforward method (e.g., full parameters) to fine-tune the model; however, the results suggest that this is less effective than the Magic Training Strategy [83]. More details and qualitative analysis are provided in Appendix D.3.
@@
Guideline for Model Selection. With the increasing number of T2V models, the community faces challenges in selecting the most appropriate model due to the tendency of each model to showcase its best results. To address this issue, we provide a guideline for model selection based on the evaluation results of ChronoMagic-Bench: (a) Except for MagicTime [83], CogVideoX [81] and Dream Machine [46], most T2V models exhibit minimal metamorphic amplitude and cannot generate complete processes rich in physical changes, such as seed germination, sunrise, or building construction; (b) The visual quality of a single frame may be high, but when viewed in sequence, flickering often occurs, indicating poor temporal coherence. This issue is particularly evident in T2V-zero [28] and OpenSora 1.2 [90], whereas closed-source models do not exhibit this problem; (c) The emergence of Sora [9] has promoted the rapid development of DiT-based methods. Closed-source models based on DiT have comprehensively surpassed those based on U-Net. However, most open-source models’ visual quality, text-following capability, and metamorphic amplitude still lag behind U-Net-based methods. We speculate that DiT-based models are more scalable and require more data, giving closed-source models a significant advantage over open-source models; (d) It is expensive to access massive data and computing resources. First, they can build datasets by crawling videos without copyright disputes. Second, adopting the U-DiT architecture may balance performance and cost to a certain extent; (e) Ordinary users who want to try T2V models can prioritize Dream Machine [46] and KeLing [33]. Researchers who wish to conduct in-depth research on T2V can prioritize the study of metamorphic video generation with CogVideoX [81], EasyAnimate [77], OpenSoraPlan [41] and OpenSora [90], as neither open-source nor closed-source models can achieve this function.
@@ 6 Conclusion
In this paper, we present ChronoMagic-Bench, the first benchmark specifically designed to assess the generation of time-lapse videos. It addresses the shortcomings of current benchmarks, which primarily focus on standard videos and overlook critical elements such as metamorphic amplitude and coherence. Additionally, we introduce two new automated metrics, MTScore and CHScore, which align with human perception. Based on ChronoMagic-Bench, we conduct a comprehensive evaluation of almost all open&closed-source leading text-to-video (T2V) models and provide crucial insights into the strengths and weaknesses of various models. Moreover, we propose ChronoMagic-Pro, the first large-scale time-lapse T2V dataset, to facilitate further research by the community.
7 Limitations and Future Work
While ChronoMagic-Bench offers a robust evaluation framework, there are still two limitations of this work. (1) The majority of the data is sourced from YouTube, where video quality varies significantly, necessitating extensive filtering based on aesthetic criteria, views, and likes. Additionally, most videos are licensed under CC BY 4.0, limiting their use to academic research. (2) Despite introducing MTScore and CHScore for metamorphic attributes and temporal coherence, a clear gap remains between these existing metrics and human preferences. Future efforts will focus on aligning automated metrics more closely with human judgment for a more accurate evaluation of T2V models.
8 Acknowledgments
We thank all the anonymous reviewers for their constructive comments. This work was supported in part by the Natural Science Foundation of China (No. 62202014, 62332002, 62425101), and Shenzhen Basic Research Program (No.JCYJ20220813151736001).
Appendix A Related Works: Text-to-Video Generation Models
@@ B.1 Construction of retrieval sentences for Metamorphic Score
B.2 Further Description of Temporal Coherence Score
@@
We present a detailed description of the algorithm for computing the Temporal Coherence Score. Specifically, we first process input video using the pre-trained model with grid size and threshold to get visibility of point . Then, we count the number of missing tracking points in each frame, and the change in missed points between consecutive frames :
where , represents the position of the frame, identifies different tracking points, and indicates the visibility of point in frame . To make the CHScore robust to temporally coherent disappearance of points, we first calculate the direction of camera/object movement based on the tracking points across all frames. Then, if the tracking point j of frame i disappears in the far direction, it is not included in the calculation of . Based on these, we then calculate the , which represents the average proportion of missed points per frame in the video. And the , which measures the variation in the number of missed points between consecutive frames, indicating frame-to-frame coherence:
where , is the mean of , is the total number of frames and is the number of points per frame. In addition, we need to calculate the , which indicates the ratio of frames that need to be cut to the total number of frames, reflecting the extent of video editing required. And the , which indicates the number of consecutive changes in missed points exceeding the threshold, indicating frequent large-scale instability in point tracking:
where represents the normalized variable, and denotes the corresponding weight coefficient. For the setting of to , we follow the following principles. Specifically, is a global metric representing the model’s overall performance across the entire video and holds the highest significance, with a weight of . measures the stability of missed points between frames, a critical aspect of video analysis, and is therefore assigned a weight of . indicates abnormal situations and carries a weight of . , similar in function to the , serves as a secondary indicator, also weighted at . Lastly, represents individual extreme cases and is assigned a lower weight of .
B.3 Details of Temporally Coherent Disappearance of Points
The ’Temporally coherent disappearance of points’ describes the phenomenon where tracking points vanish over time due to movements such as camera movement or water flow, potentially causing these points to exit the camera’s field of view. To prevent this from influencing the CHScore calculation, we initially identify the direction of change for various points within the video, as depicted in Figure 7. Subsequently, points that vanish proximate to this directional endpoint are excluded from the CHScore calculation.
B.4 Visual Reference of the Different Scores of MTScore and CHScore